# RealRAG: Retrieval-augmented Realistic Image Generation via Self-reflective Contrastive Learning

### v1

@@ 1 Introduction
Recent text-to-image generators have achieved notable progress in image synthesis from the given textual prompts. There are three mainstream types of generative models, including the U-Net-based diffusion model (Rombach et al., 2022; Podell et al., 2023), the DiT-based diffusion model (Xiao et al., 2024; Sun et al., 2024), and the auto-regressive model (Esser et al., 2024; BlackForest, 2024). Typically, these models store all their visual memory (e.g., the appearance of Big Ben) implicitly in the parameters of the underlying neural network, requiring a lot of parameters(e.g., 10B). Furthermore, similar to the hallucination problem of Large Language Models (LLMs) (OpenAI, 2023; Touvron et al., 2023), the large-scale text-to-image generative models also show the same problem. Some generated images include ghosting, distortions, and unnatural elements when generating specific real-world objects. Therefore, these problems motivate the development of text-to-image generation models, which can integrate external visual knowledge (e.g., images from the web) to augment generative realism and accuracy for fine-grained and unseen novel object generation.
@@
We apply our RealRAG to all types of state-of-the-art text-to-image generative models, including U-Net-based diffusion models (SD V2.1 (Rombach et al., 2022), SD XL (Podell et al., 2023)), DiT-based diffusion models (SD V3 (Esser et al., 2024), Flux (BlackForest, 2024)), and auto-regressive models (OmniGen (Xiao et al., 2024), Emu (Sun et al., 2024)). Note that, our RealRAG is the first work to build a unified RAG framework for all types of text-to-image generative models. Our RealRAG consistently delivers significant performance improvements with all these models, such as +5.77% on the Stanford Cars benchmark with Flux (BlackForest, 2024) and +20.48% on the Oxford Flowers benchmark with Emu (Sun et al., 2024). Additionally, to evaluate the capability of unseen novel object generation, we collect recent novel objects from news web pages and construct human evaluations. Our RealRAG demonstrates a substantial performance improvement in human evaluation. Code will be available upon acceptance.
@@ 2.1 Text-to-Image Generation
The U-Net-based Stable Diffusion models (Rombach et al., 2022; Podell et al., 2023) first perform universal image generation from text prompts, which typically trained on large scale text-image paired dataset, a.k.a., LIANG 5B (Schuhmann et al., 2022). After the proposal of the Diffusion Transformer (DiT) (Peebles & Xie, 2023), some research, such as Stable Diffusion V3 (Esser et al., 2024) and Flux (BlackForest, 2024), utilize DiT as the backbone to develop DiT-based Diffusion Models for text-to-image generation. Recently, with the success of auto-regressive (AR) modeling in the natural language processing (OpenAI, 2023; Touvron et al., 2023), some works have explored how to combine auto-regressive models with diffusion models to improve the understanding capability and further build a unified multi-modal model for both understanding and generation (Chen et al., 2024; Xie et al., 2024; Zhou et al., 2024). These AR-based models, such as OmniGen (Xiao et al., 2024) and Emu (Sun et al., 2024), also show notable performance on the text-to-image task. While these methods have achieved strong performance in text-to-image generation, they store all the knowledge in their pre-trained parameters, which leads to hallucinations and distortions when generating realistic objects. To address this limitation, we propose the real-object-based RAG framework to integrate missing knowledge and improve the ability to generate realistic images.
@@ 2.2 Retrieval-augmented Generation
Retrieval-augmented generation has shown promise with NLP (Lewis et al., 2020; Guu et al., 2020). To incorporate external knowledge into a LLM (Gao et al., 2023; Jiang et al., 2023), these methods retrieve documents relevant to inputs from an external database, subsequently, the LLM utilizes the recalled documents as references to generate accurate results. The external knowledge used is typically a text database (Hashimoto et al., 2018; Khandelwal et al., 2019; Shi et al., 2023). However, the text database is not direct and controllable for realistic image generation (Blattmann et al., 2022). In this paper, we conduct a vision-based, real-object-based database, which is collected by realistic images from public real-world datasets, including ImageNet (Deng et al., 2009), Stanford Cars (Krause et al., 2013), Stanford Dogs (Dataset, 2011), and Oxford Flowers (Nilsback & Zisserman, 2008). In this way, we augment the realism of the generative images with the real-object-based database.
@@ 3.1 Self-reflective Contrastive Learning
$\mathcal{D}_{neg}^{nor}=\sum_{j\neq pos}^{N}(Z_{q}^{T}\cdot Z_{obj}^{j}/\tau),\quad\mathcal{D}_{neg}^{ref}=exp(Z_{q}^{T}\cdot Z_{neg}),$ (7)
@@
$\mathcal{L}=-log\frac{exp(Z_{q}^{T}\cdot Z_{obj}^{pos})/\tau}{exp(Z_{q}^{T}\cdot Z_{obj}^{pos}/\tau)+\mathcal{D}_{neg}^{nor}+\mathcal{D}_{neg}^{ref}},$ (8)
@@ 3.2 Real-object-based Retrieval-augmented Generation
[…]
SD V2.1!(Rombach et al., 2022) 61.48 14.16 64.99 68.80 17.69 49.60 82.97 20.76 69.17
[…]
@@
[…]
SD V2.1!(Rombach et al., 2022) 37.45
[…]
@@ 3.3 Implementation
Generator. We use existing text-to-image generative models as the generator to implement our RealRAG. Concretely, we implement our RealRAG with the following models: U-Net-based diffusion models (SD V2.1 (Rombach et al., 2022), SD XL (Podell et al., 2023)), DiT-based diffusion models (SD V3 (Esser et al., 2024), Flux (BlackForest, 2024)), and auto-regressive models (OmniGen (Xiao et al., 2024), Emu (Sun et al., 2024)).
@@ 4.2.2 Generative Result
Qualitative Results. We show the qualitative results in Fig. 4. The visual results show the ability of our RealRAG to overcome hallucinations and significantly improve the realism and quality of fine-grained realistic image generation. Specifically, for the autoregressive model, In the prompt [”Audi 100 Wagon 1994”], the RealRAG model produces a highly realistic rendering of the vehicle, capturing intricate details like the side mirrors and headlights, which are closer to the real object compared to the original AR model with significant hallucination in the car’s side. For the U-Net-based diffusion model, In prompt [”Silverbush”], RealRAG ensures the correct flower structure and vibrant color tones, whereas the original model struggles with maintaining fine-grained details like petal shape. Lastly, compared with the original generator, RealRAG delivers a compelling representation of the breed’s signature coat patterns and body posture from the [”Bluetick”] prompt. The results demonstrate that RealRAG generates high realism and correct objects and show the flexible application of the SoTA generators.
@@ 4.3.1 Experiment Setup
To evaluate the ability to generate unseen novel objects, we collect several recently introduced objects (e.g., the Cybertruck) to form the input prompts. Specifically, we use a LLM (e.g., ChatGPT) to generate prompts based on these novel objects. To ensure fairness in our comparison, we conduct both qualitative evaluations and human evaluations.
@@ A.1 More Experimental Results
[…]
SD V2.1(Rombach et al., 2022) 35.21 42.35 34.79 37.45
[…]

### v2

@@ 1 Introduction
Recent text-to-image generators have achieved notable progress in image synthesis from the given textual prompts. There are three mainstream types of generative models, including the U-Net-based diffusion model (Rombach et al., 2022a; Podell et al., 2023), the DiT-based diffusion model (Xiao et al., 2024; Sun et al., 2024), and the auto-regressive model (Esser et al., 2024; BlackForest, 2024). Typically, these models store all their visual memory (e.g., the appearance of Big Ben) implicitly in the parameters of the underlying neural network, requiring a lot of parameters(e.g., 10B). Furthermore, similar to the hallucination problem of Large Language Models (LLMs) (OpenAI, 2023; Touvron et al., 2023), the large-scale text-to-image generative models also show the same problem. Some generated images include ghosting, distortions, and unnatural elements when generating specific real-world objects. Therefore, these problems motivate the development of text-to-image generation models, which can integrate external visual knowledge (e.g., images from the web) to augment generative realism and accuracy for fine-grained and unseen novel object generation.
@@
We apply our RealRAG to all types of state-of-the-art text-to-image generative models, including U-Net-based diffusion models (SD V2.1 (Rombach et al., 2022a), SD XL (Podell et al., 2023)), DiT-based diffusion models (SD V3 (Esser et al., 2024), Flux (BlackForest, 2024)), and auto-regressive models (OmniGen (Xiao et al., 2024), Emu (Sun et al., 2024)). Note that, our RealRAG is the first work to build a unified RAG framework for all types of text-to-image generative models. Our RealRAG consistently delivers significant performance improvements with all these models, such as +5.77% on the Stanford Cars benchmark with Flux (BlackForest, 2024) and +20.48% on the Oxford Flowers benchmark with Emu (Sun et al., 2024). Additionally, to evaluate the capability of unseen novel object generation, we collect recent novel objects from news web pages and construct human evaluations. Project page: https://qc-ly.github.io/RealRAG-page/
@@ 2.1 Text-to-Image Generation
The U-Net-based Stable Diffusion models (Rombach et al., 2022a; Podell et al., 2023) first perform universal image generation from text prompts, which typically trained on large scale text-image paired dataset, a.k.a., LAION 5B (Schuhmann et al., 2022). After the proposal of the Diffusion Transformer (DiT) (Peebles & Xie, 2023), some research, such as Stable Diffusion V3 (Esser et al., 2024) and Flux (BlackForest, 2024; Liu et al., 2025), utilize DiT as the backbone to develop DiT-based Diffusion Models for text-to-image generation. Recently, with the success of auto-regressive (AR) modeling in natural language processing (OpenAI, 2023; Touvron et al., 2023), some works have explored how to combine auto-regressive models with diffusion models to improve the understanding capability and further build a unified multi-modal model for both understanding and generation (Chen et al., 2024; Xie et al., 2024; Zhou et al., 2024). These AR-based models, such as OmniGen (Xiao et al., 2024) and Emu (Sun et al., 2024), also show notable performance on the text-to-image task. While these methods have achieved strong performance in text-to-image generation, they store all the knowledge in their pre-trained parameters, which leads to hallucinations and distortions when generating realistic objects. To address this limitation, we propose the real-object-based RAG framework to integrate missing knowledge and improve the ability to generate realistic images.
@@ 2.2 Retrieval-augmented Generation
Retrieval-augmented generation has shown promise with NLP (Lewis et al., 2020; Guu et al., 2020). To incorporate external knowledge into a LLM (Gao et al., 2023; Jiang et al., 2023b), these methods retrieve documents relevant to inputs from an external database, subsequently, the LLM utilizes the recalled documents as references to generate accurate results. The external knowledge used is typically a text database (Hashimoto et al., 2018; Khandelwal et al., 2019; Shi et al., 2023; Lyu et al., 2024b). However, the text database is not direct and controllable for realistic image generation (Blattmann et al., 2022; Zheng et al., 2025). In this paper, we conduct a vision-based, real-object-based database, which is collected by realistic images from public real-world datasets, including ImageNet (Deng et al., 2009), Stanford Cars (Krause et al., 2013), Stanford Dogs (Dataset, 2011), and Oxford Flowers (Nilsback & Zisserman, 2008). In this way, we augment the realism of the generative images with the real-object-based database.
@@ 3.1 Self-reflective Contrastive Learning
$\mathcal{D}_{neg}^{nor}=\sum_{j\neq pos}^{N}(exp(Z_{q}^{T}\cdot Z_{obj}^{j}/\tau)),\quad\mathcal{D}_{neg}^{ref}=exp(Z_{q}^{T}\cdot Z_{neg}),$ (7)
@@
$\mathcal{L}=-log\frac{exp(Z_{q}^{T}\cdot Z_{obj}^{pos}/\tau)}{exp(Z_{q}^{T}\cdot Z_{obj}^{pos}/\tau)+\mathcal{D}_{neg}^{nor}+\mathcal{D}_{neg}^{ref}},$ (8)
@@ 3.2 Real-object-based Retrieval-augmented Generation
Emu W. Ours 63.07 15.30 70.55 80.39 19.60 45.04 85.52 22.42 63.85
$\Delta$ +1.54 +2.52 -9.28 +3.08 +0.23 -4.28 +2.76 +0.75 -16.14
SD V2.1(Rombach et al., 2022a) 61.48 14.16 64.99 68.80 17.69 49.60 82.97 20.76 69.17
U-Net-based SD V2.1 W. Ours 64.50 15.84 58.98 79.06 18.57 43.93 88.43 21.77 60.76
@@
Emu W. Ours 39.90
$\Delta$ +3.89
SD V2.1(Rombach et al., 2022a) 37.45
U-Net-based SD V2.1 W. Ours 40.23
@@ 3.3 Implementation
Generator. We use existing text-to-image generative models as the generator to implement our RealRAG. Concretely, we implement our RealRAG with the following models: U-Net-based diffusion models (SD V2.1 (Rombach et al., 2022a), SD XL (Podell et al., 2023)), DiT-based diffusion models (SD V3 (Esser et al., 2024), Flux (BlackForest, 2024)), and auto-regressive models (OmniGen (Xiao et al., 2024), Emu (Sun et al., 2024)). Specifically, for autoregressive models, they can directly generate images from the image condition. For diffusion models, we utilize ControlNet, which introduces a branch to stable diffusion, enabling the inclusion of additional inputs, to input image-based conditions. For example. First, we retrieve and sort the closest images. Next, we input the selected images into the ControlNet branch to control specific elements during the image synthesis process.
@@ 4.2.2 Generative Result
Qualitative Results. We show the qualitative results in Fig. 4. The visual results show the ability of our RealRAG to overcome hallucinations and significantly improve the realism and quality of fine-grained realistic image generation. Specifically, for the autoregressive model, In the prompt [”Audi TT Hatchback 2011”], the RealRAG model produces a highly realistic rendering of the vehicle, capturing intricate details like the side mirrors and headlights, which are closer to the real object compared to the original AR model with significant hallucination in the car’s side. For the U-Net-based diffusion model, In prompt [”Silverbush”], RealRAG ensures the correct flower structure and vibrant color tones, whereas the original model struggles with maintaining fine-grained details like petal shape. Lastly, compared with the original generator, RealRAG delivers a compelling representation of the breed’s signature coat patterns and body posture from the [”Bluetick”] prompt. The results demonstrate that RealRAG generates high realism and correct objects and show the flexible application of the SoTA generators.
@@ 4.3.1 Experiment Setup
To evaluate the ability to generate unseen novel objects, we collect several recently introduced objects (e.g., the Cybertruck) to form the input prompts. Specifically, we use a LLM (e.g., ChatGPT) to generate prompts based on these novel objects. For the image database, we use the images from the Internet (Google Images), which include the novel objects. To ensure fairness in our comparison, we conduct both qualitative evaluations and human evaluations.
@@ 5 Ablation Study
6 Discussion
Research Purpose: Different from RDMs (Rombach et al., 2022b; Blattmann et al., 2022; Sheynin et al., 2022; Chen et al., 2022) that use retrieval-augmented techniques to train or fine-tune a diffusion model and achieve out-of-distribution (OOD) image generation by switching databases, our RealRAG aims to use retrieval-augmented techniques to train or fine-tune a diffusion model and achieve out-of-distribution (OOD) image generation by switching databases.
Unseen novel objects: These refer to objects that appear after the generative models and retrieval models are trained. The generative model cannot generate these objects, and the retrieval model can’t easily retrieve relevant references by similarity. This is a much more challenging issue.
Hallucination When Generating Fine-Grained Objects: Existing SoTA t2i models are pre-trained on large-scale text-image paired datasets. As a result, they tend to produce hallucinations, such as inaccuracy or unrealistic features, when generating fine-grained objects. This is a problem inherent to large generative models.
@@ 7 Conclusion
Acknowledgments
This work was supported by Open Project Program of Guangxi Key Laboratory of Digital Infrastructure (Grant Number: GXDIOP2024015); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ A.1 More Experimental Results
Emu W. Ours 40.24 44.33 35.13 39.90
$\Delta$ +5.48 +3.40 +2.82 +3.89
SD V2.1(Rombach et al., 2022a) 35.21 42.35 34.79 37.45
U-Net-based SD V2.1 W. Ours 39.43 46.10 35.15 40.23
@@ B.2 Generative Results
Appendix C Application Value
For existing commercial text-to-image generative models, training them is cost-prohibitive. Therefore, a pipeline is needed to integrate real-time updated data from the internet into the generative model without any retraining, enabling the generation of unseen novel objects. On the other hand, in specific application scenarios such as advertising creation and multi-modal generation (Tang et al., 2023; Lyu et al., 2024a), users need generated images that meet their design requirements, while also ensuring that products (fine-grained objects) within the image remain realistic. This requires generative models to have the ability to generate both open-domain and fine-grained objects. Therefore, RealRAG focuses on reducing hallucinations in large t2i generators through RAG technology, enabling open-domain generators to generate specific fine-grained objects.
Beyond text-to-image (T2I) generation itself, several downstream tasks depend on and extend the same technology, like 3D generation (Li et al., 2024; Jiang et al., 2023a; Li et al., 2023) and video generation (Singer et al., 2022; Gu et al., 2025; Wang et al., 2025). For example, (i) text-to-3D synthesis (Poole et al., 2022; Jiang et al., 2024a, b; Bai et al., 2023) generally requires a strong T2I backbone, while (ii) image-to-3D reconstruction (Jiang et al., 2025; Hong et al., 2023; Tang et al., 2024) often begins with an image created by a T2I model.